# A Theory of the NEPv Approach for Optimization On the Stiefel Manifold

### v2

@@ 8.2 Not all $P_{i}$ are the entire $P$
$\operatorname{tr}(P_{i}^{\operatorname{T}}H_{i}(P_{i})P_{i})$ $=\underaccent{\bar}{\gamma}_{i}\,f_{i}(P_{i})\quad\mbox{for $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$},$ (8.16a)
and given $\widehat{P}\in\mathbb{O}^{n\times k}$ and $P\in\mathbb{P}$, there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
We seek an alternative route, as our second attempt, by resorting the generic $H(P)$ in (6.5):
$H(P)=\sum_{i=1}^{N}\phi_{i}(T(P))\left(\frac{\partial f_{i}(P_{i})}{\partial P_{i}}P_{i}^{\operatorname{T}}+P_{i}\left[\frac{\partial f_{i}(P_{i})}{\partial P_{i}}\right]^{\operatorname{T}}\right).$ (8.17)
In doing so, we completely ignore $H_{i}(P_{i})$ for $1\leq i\leq N$ that we presumably already know and should take advantage of. To proceed, we will also need the same consistency conditions as in (5.5) for NPDo. This makes all components of $T(P)$ atomic functions for NPDo and for NEPv. Hence the NPDo approach can be applied in the first place. Nonetheless, the NEPv approach will also work with $H(P)$, assuming (5.5).
@@
Consider $f=\phi\circ T$, where $T(\cdot)$ takes the form in (5.3) and $\phi$ is convex and differentiable with partial derivatives denoted by $\phi_{i}$ as in (5.2). Let $H(P)$ be given by (8.17). Suppose that the conditions in (5.5) hold and that $\phi_{i}(\boldsymbol{x})\geq 0$ for those $i$ for which (5.5b) does not become an equality. Then the NEPv Ansatz holds with $\omega=1/(2\alpha)$.
@@
Given $\widehat{P}\in\mathbb{O}^{n\times k}$ and $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$, suppose that (6.7) holds, i.e., $\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})\geq\operatorname{tr}(P^{\operatorname{T}}H(P)P)+\eta$. Write
@@
where $x_{i}=f_{i}(P_{i})$ and $\widetilde{x}_{i}=f_{i}(\widetilde{P}_{i})$, and write $\mathscr{H}_{i}(P_{i})=\frac{\partial f_{i}(P_{i})}{\partial P_{i}}$. For $H(P)$ in (8.17), we have by (5.5)
$\operatorname{tr}(P^{\operatorname{T}}H(P)P)$ $=2\sum_{i=1}^{N}\phi_{i}(\boldsymbol{x})\operatorname{tr}(P_{i}^{\operatorname{T}}\mathscr{H}_{i}(P_{i}))$
@@
Let $W_{1}\in\mathbb{O}^{k\times k}$ be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}\mathscr{H}(P)$ and $Z=\widehat{P}W_{1}$. We have
$\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})$ $=2\sum_{i=1}^{N}\phi_{i}(\boldsymbol{x})\operatorname{tr}(\widehat{P}^{\operatorname{T}}\mathscr{H}_{i}(P_{i})P_{i}^{\operatorname{T}}\widehat{P})$
$=2\sum_{i=1}^{N}\phi_{i}(\boldsymbol{x})\operatorname{tr}(\widehat{P}^{\operatorname{T}}\mathscr{H}_{i}(P_{i})J_{i}^{\operatorname{T}}P^{\operatorname{T}}\widehat{P})$
$=2\operatorname{tr}(\widehat{P}^{\operatorname{T}}\mathscr{H}(P)P^{\operatorname{T}}\widehat{P})$
$\leq 2\|\widehat{P}^{\operatorname{T}}\mathscr{H}(P)\|_{\operatorname{tr}}$
$=2\operatorname{tr}(Z^{\operatorname{T}}\mathscr{H}(P))$
$=2\sum_{i=1}^{N}\phi_{i}(\boldsymbol{x})\operatorname{tr}(Z_{i}^{\operatorname{T}}\mathscr{H}_{i}(P_{i})),$
where $Z_{i}=ZJ_{i}$ for $1\leq i\leq N$ are submatrices of $Z$. Now use the second consistency condition in (5.5b) to get $W_{2}\in\mathbb{O}^{k\times k}$ and set $\widetilde{P}=ZW_{2}=\widehat{P}(W_{1}W_{2})$. Then
$\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})\leq 2\sum_{i=1}^{N}\phi_{i}(\boldsymbol{x})\operatorname{tr}(Z_{i}^{\operatorname{T}}\mathscr{H}_{i}(P_{i}))\leq 2\sum_{i=1}^{N}\phi_{i}(\boldsymbol{x})\,(\alpha\widetilde{x}_{i}+\beta_{i}x_{i}).$
Plug them into $\eta+\operatorname{tr}(P^{\operatorname{T}}H(P)P)\leq\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})$ and simplify the resulting inequality with the help of $\gamma_{i}=\alpha+\beta_{i}$ to get
@@
We now explain why the resulting Algorithm 1 is not going to be competitive to Algorithm 1 per SCF iterative step. First both algorithms work due to the same set of consistency conditions in (5.5). Next carefully examining the proof of Theorem 8.3, we find that $Q_{i}$ at Line 4 of Algorithm 1 is the product of two $k\times k$ orthogonal matrices, $W_{1}$ from an orthogonal polar factor of $[\widehat{P}^{(i)}]^{\operatorname{T}}\mathscr{H}(P^{(i)})$ and $W_{2}$ from the second consistency condition in (5.5b). Algorithm 1 also needs $Q_{i}$ at its line 4 but just one orthogonal matrix dictated by the second consistency condition in (5.5), making its SCF step cheaper, not to mention there is an additional partial eigendecomposition to compute at Line 3 of Algorithm 1.
@@
which is positive semidefinite for $x>0$ and $y\geq 0$ if and only if $20\leq\theta\leq 1/$ 2. This implies $\phi(\boldsymbol{x})\equiv\phi_{0}(x_{1},x_{2}+x_{3})$ is convex for $x_{1}>0$ and $x_{2}+x_{3}\geq 0$, provided $20\leq\theta\leq 1/$ 2. However, the results in [67] says that the NEPv approach works on $f$ for $\Theta$TR for $0\leq\theta\leq 1$, much bigger range for $\theta$ than what Theorem 8.2 for $f^{2}$ implies. Theorem 8.2 also yields an inequality on how much the objective function squared, $f^{2}$, increases, in contrast to the previous (6.11) which is for the original objective function, $f$, for $0\leq\theta\leq 1$ and OCCA. In any case, the best $Q_{i}$ at Line 4 of Algorithm 1 when applied to $\Theta$TR is an orthogonal polar factor of $[\widehat{P}^{(i)}]^{\operatorname{T}}D$ and $\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$.
@@ 9 A brief comparison of the NPDo and NEPv approaches
The NEPv approach requires weaker conditions.
We have demonstrated that the NDPo Ansatz demands more on an objective function than the NEPv Ansatz, and hence the NEPv approach provably works on a wider collection of maximization problems (1.1) on the Stiefel manifold than the NPDo approach.
@@
    As a further demonstration, in Table 6, we summarize what are required by both approaches on three convex compositions of matrix-trace functions, and it clearly indicates that NEPv requires weaker conditions than NPDo does for both $\phi\circ T_{1a}$ and $\phi\circ T_{2a}$.
The NPDo approach is easier to use.
The NPDo approach, if it provably works, is easier to implement and more flexible to use.
@@ 10 Concluding Remarks
This is an $n\times k$ matrix equation in $P$ on the Stiefel manifold, upon noticing $\Lambda=[P^{\operatorname{T}}\mathscr{H}(P)+\mathscr{H}(P)^{\operatorname{T}}P]/2$. Any maximizer is a solution. Except for very special objective functions such as $\operatorname{tr}(P^{\operatorname{T}}AP)$, solving this equation rightly and efficiently for a maximizer is a challenging task. For example, it often has infinitely many solutions and maximizers hide among them. Hence we need to not only solve the nonlinear equation of the KKT condition but also find the right ones. Inspired by recent works [67, 66, 74, 75, 76, 77], in this paper, we establish two unifying frameworks, one for the NEPv approach and one for the NPDo approach, for optimization on the Stiefel manifold. Our frameworks are built upon two fundamental ansatzes, the NPDo Ansatz and the NEPv Ansatz. When a respective ansatz is satisfied, the corresponding approach, the NPDo or NEPv approach, is guaranteed to work in the sense of global convergence from any given initial point. To expand the applicability of the approaches, we propose the theories of atomic functions for each approach and show that any convex composition $\phi\circ T$ of atomic functions for any of the two approaches satisfies the corresponding ansatz under some mild conditions. It is demonstrated that the commonly used matrix-trace functions
@@
However when an ansatz fails to hold for a given objective function, the conclusions from our global convergence analysis will likely fail. In the case for the NEPv approach, a remedy via the level-shifting SCF exists to ensure locally linear convergence when $f(P)$ is right-unitarily invariant [4] or when $f(P)$ contains and increases with $\operatorname{tr}(P^{\operatorname{T}}D)$ [46], where sharp estimations of linear convergence rate are obtained. But there are more work to do, especially for the NPDo approach for which a remedy remains to be found. In [62], it is investigated how $\operatorname{tr}(P^{\operatorname{T}}D)$ may help determine a particular $P$ among all orthonormal basis matrices of a subspace.
Not all objective functions (or their simple transformations like $f^{2}$ for $\Theta$TR as explained in Remark 8.1) that satisfy the ansatz(es) take the form of some convex compositions of atomic functions, even for those in Table 1. For example, in the case of OLDA, $f(P)=\phi\circ T(P)$ where $\phi(\boldsymbol{x})=x_{1}/x_{2}$ and $T(P)=\begin{bmatrix}\operatorname{tr}(P^{\operatorname{T}}AP)\\ \operatorname{tr}(P^{\operatorname{T}}BP)\end{bmatrix}$. Although this $f(P)$ is not a convex composition of $\operatorname{tr}(P^{\operatorname{T}}AP)$ and $\operatorname{tr}(P^{\operatorname{T}}BP)$, it still satisfies the NEPv Ansatz. In fact, OLDA is a special case of $\Theta$TR: $\theta=1$ and $D=0$, and inequality (6.11) applies and yields, for OLDA,

### v4

@@ 8.2 Not all $P_{i}$ are the entire $P$
$\operatorname{tr}(P_{i}^{\operatorname{T}}H_{i}(P_{i})P_{i})$ $=\underaccent{\bar}{\gamma}_{i}\,f_{i}(P_{i})\quad\mbox{for $P\in\mathbb{P}\subseteq{\rm St}(k,n)$},$ (8.16a)
and given $\widehat{P}\in{\rm St}(k,n)$ and $P\in\mathbb{P}$, there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
We will seek an alternative route, as our second attempt. Assume the following setting: among the $N$ components $f_{i}(P_{i})$, the first $N_{1}$ of them involve $P_{i}$ that are not the entire $P$ but the last $N_{2}$ do, i.e.,
$P_{i}=P\quad\mbox{for $N_{1}+1\leq i\leq N$},$ (8.17)
where $N_{1}+N_{2}=N$. We resort the generic symmetric matrix-valued function in (6.5) for the first $N_{1}$ components $f_{i}(P_{i})$ for $1\leq i\leq N_{1}$ but the individual $H_{i}(P)$ for the last $N_{2}$ components $f_{i}(P)$ for $N_{1}+1\leq i\leq N$ to get:
$H(P)=\sum_{i=1}^{N_{1}}\phi_{i}(T(P))\left(\frac{\partial f_{i}(P_{i})}{\partial P_{i}}P_{i}^{\operatorname{T}}+P_{i}\left[\frac{\partial f_{i}(P_{i})}{\partial P_{i}}\right]^{\operatorname{T}}\right)+\sum_{i=N_{1}+1}^{N}\phi_{i}(T(P))\,H_{i}(P).$ (8.18)
In doing so, we completely ignore $H_{i}(P_{i})$ for $1\leq i\leq N_{1}$ that we presumably already know and should take advantage of but cannot. To proceed, we will also need the same consistency conditions as in (5.5) for NPDo for $f_{i}(P_{i})$ for $1\leq i\leq N_{1}$ but the ones in (8.16) for $f_{i}(P)$ for $N_{1}+1\leq i\leq N$. This makes the first $N_{1}$ components of $T(P)$ atomic functions for both NPDo and for NEPv.
In this setting $N_{1}=0$ or $N_{2}=0$ are allowed. The case $N_{1}=0$ is the one we have already dealt with in subsection 8.1, and for the case $N_{2}=0$, the NPDo approach can be applied in the first place. It remains to conquer the case both $N_{1},\,N_{2}\geq 1$. Our next theorem will help us to do that.
@@
Consider $f=\phi\circ T$, where $T(\cdot)$ takes the form in (5.3) with (8.17) and $\phi$ is convex and differentiable with partial derivatives denoted by $\phi_{i}$ as in (5.2). Let $H(P)$ be given by (8.18). Suppose that
  1. (i)
    the conditions in (5.5) hold for $1\leq i\leq N_{1}$,
  2. (ii)
    the conditions in (8.16) hold with $Q=I_{k}$ for $N_{1}+1\leq i\leq N$,
  3. (iii)
    $\phi_{i}(\boldsymbol{x})\geq 0$ for those $1\leq i\leq N_{1}$ for which (5.5b) does not become an equality and for those $N_{1}+1\leq i\leq N$ for which (8.16b) does not become an equality,
  4. (iv)
    $\underaccent{\bar}{\alpha}=2\alpha$ for $\alpha$ in (5.5b) and $\underaccent{\bar}{\alpha}$ in (8.16b).
Then the NEPv Ansatz holds with $\omega=1/(2\alpha)$ and with the $Q$-matrix specified in the proof.
@@
Given $\widehat{P}\in{\rm St}(k,n)$ and $P\in\mathbb{P}\subseteq{\rm St}(k,n)$, suppose that (6.7) holds, i.e., $\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})\geq\operatorname{tr}(P^{\operatorname{T}}H(P)P)+\eta$. Write
@@
where $x_{i}=f_{i}(P_{i})$ and $\widetilde{x}_{i}=f_{i}(\widetilde{P}_{i})$, and write $\mathscr{H}_{i}(P_{i})=\frac{\partial f_{i}(P_{i})}{\partial P_{i}}$. For $H(P)$ in (8.18), we have by (5.5) and (8.16)
$\operatorname{tr}(P^{\operatorname{T}}H(P)P)$ $=2\sum_{i=1}^{N_{1}}\phi_{i}(\boldsymbol{x})\operatorname{tr}(P_{i}^{\operatorname{T}}\mathscr{H}_{i}(P_{i}))+\sum_{i=N_{1}+1}^{N}\phi_{i}(\boldsymbol{x})\operatorname{tr}(P^{\operatorname{T}}H_{i}(P)P)$
@@
[…]
We have
$\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})$ $=2\sum_{i=1}^{N_{1}}\phi_{i}(\boldsymbol{x})\operatorname{tr}(\widehat{P}^{\operatorname{T}}\mathscr{H}_{i}(P_{i})P_{i}^{\operatorname{T}}\widehat{P})+\sum_{i=N_{1}+1}^{N}\phi_{i}(\boldsymbol{x})\,\operatorname{tr}(\widehat{P}^{\operatorname{T}}H_{i}(P)\widehat{P})$
$=2\sum_{i=1}^{N_{1}}\phi_{i}(\boldsymbol{x})\operatorname{tr}(\widehat{P}^{\operatorname{T}}\mathscr{H}_{i}(P_{i})J_{i}^{\operatorname{T}}P^{\operatorname{T}}\widehat{P})+\sum_{i=N_{1}+1}^{N}\phi_{i}(\boldsymbol{x})\,\operatorname{tr}(\widehat{P}^{\operatorname{T}}H_{i}(P)\widehat{P})$
[…]
$=2\sum_{i=1}^{N_{1}}\phi_{i}(\boldsymbol{x})\operatorname{tr}(Z_{i}^{\operatorname{T}}\mathscr{H}_{i}(P_{i}))+\sum_{i=N_{1}+1}^{N}\phi_{i}(\boldsymbol{x})\,\operatorname{tr}(Z^{\operatorname{T}}H_{i}(P)Z),$
where $Z_{i}=ZJ_{i}$ for $1\leq i\leq N$ are submatrices of $Z$. Now use the second consistency condition in (5.5b) to get $W_{2}\in{\rm St}(k,k)$ and set $\widetilde{P}=ZW_{2}=\widehat{P}(W_{1}W_{2})=:\widehat{P}Q$. Then
$\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})$ $\leq 2\sum_{i=1}^{N_{1}}\phi_{i}(\boldsymbol{x})\operatorname{tr}(Z_{i}^{\operatorname{T}}\mathscr{H}_{i}(P_{i}))+\sum_{i=N_{1}+1}^{N}\phi_{i}(\boldsymbol{x})\,\operatorname{tr}(\widetilde{P}^{\operatorname{T}}H_{i}(P)\widetilde{P})$
$\leq 2\sum_{i=1}^{N_{1}}\phi_{i}(\boldsymbol{x})\,(\alpha\widetilde{x}_{i}+\beta_{i}x_{i})+\sum_{i=N_{1}+1}^{N}\phi_{i}(\boldsymbol{x})\,(\underaccent{\bar}{\alpha}\widetilde{x}_{i}+\underaccent{\bar}{\beta}_{i}x_{i})$
$\leq 2\sum_{i=1}^{N_{1}}\phi_{i}(\boldsymbol{x})\,(\alpha\widetilde{x}_{i}+\beta_{i}x_{i})+\sum_{i=N_{1}+1}^{N}\phi_{i}(\boldsymbol{x})\,(2\alpha\widetilde{x}_{i}+\underaccent{\bar}{\beta}_{i}x_{i}).$
Plug them into $\eta+\operatorname{tr}(P^{\operatorname{T}}H(P)P)\leq\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})$ and simplify the resulting inequality with the help of $\gamma_{i}=\alpha+\beta_{i}$ for $1\leq i\leq N_{1}$ and $\underaccent{\bar}{\gamma}_{i}=2\alpha+\underaccent{\bar}{\beta}_{i}$ for $N_{1}+1\leq i\leq N$ to get
@@
We now explain why the resulting Algorithm 1 is not going to be competitive to Algorithm 1 per SCF iterative step in the case $N_{2}=0$. First both algorithms work due to the same set of consistency conditions in (5.5). Next carefully examining the proof of Theorem 8.3, we find that $Q_{i}$ at Line 4 of Algorithm 1 is the product of two $k\times k$ orthogonal matrices, $W_{1}$ from an orthogonal polar factor of $[\widehat{P}^{(i)}]^{\operatorname{T}}\mathscr{H}(P^{(i)})$ and $W_{2}$ from the second consistency condition in (5.5b). Algorithm 1 also needs $Q_{i}$ at its line 4 but just one orthogonal matrix dictated by the second consistency condition in (5.5), making its SCF step cheaper, not to mention there is an additional partial eigendecomposition to compute at Line 3 of Algorithm 1. Having said that, the significance of Theorem 8.3 lies in the mixture case: both $N_{1}\geq 1$ and $N_{2}\geq 1$, for which the theorem ensures that Algorithm 1 can still be applied with guaranteed convergence. The next example falls into such a category.
Example 8.4.
Consider $T_{4}$ that shares some similarity with $T_{2}$ in (5.10) and $T_{2a}$ in (8.8):
$T_{4}\,:\,P\in{\rm St}(k,n)\to T_{4}(P):=\begin{bmatrix}\|P_{1}^{\operatorname{T}}A_{1}P_{1}\|_{\operatorname{F}}^{2}\\ \vdots\\ \|P_{\ell}^{\operatorname{T}}A_{\ell}P_{\ell}\|_{\operatorname{F}}^{2}\\ \operatorname{tr}(P^{\operatorname{T}}B_{1}P)\\ \vdots\\ \operatorname{tr}(P^{\operatorname{T}}B_{t}P)\end{bmatrix}\in\mathbb{R}^{\ell+t},$ (8.19)
where $A_{i}\succeq 0$ for $1\leq i\leq\ell$ and $B_{i}\in\mathbb{R}^{n\times n}$ for $1\leq i\leq t$ are symmetric. Each $\|P_{i}^{\operatorname{T}}A_{i}P_{i}\|_{\operatorname{F}}^{2}=\operatorname{tr}\big{(}\big{[}P_{i}^{\operatorname{T}}A_{i}P_{i}\big{]}^{2}\big{)}$ is an atomic function of NPDo by Theorem 4.4, and each $\operatorname{tr}(P^{\operatorname{T}}B_{i}P)$ is an atomic function of NEPv (by Theorem 7.5) but may not be an atomic function of NPDo unless $B_{i}\succeq 0$ also. We can be do away with $B_{i}\not\succeq 0$ by shifting $B_{i}$ to $B_{i}-\delta I\succeq 0$ as we did in the second part of Example 8.2, but the shift works only if the corresponding partial derivative $\phi_{\ell+i}\geq 0$. If we do not have that, then the NPDo approach is not guaranteed to work even with the shifting technique. With Theorem 8.3, however, we can make the NEPv approach work with
$H(P)$ $=\sum_{i=1}^{\ell}\phi_{i}(T(P))\left(\frac{\partial\operatorname{tr}\big{(}\big{[}P_{i}^{\operatorname{T}}A_{i}P_{i}\big{]}^{2}\big{)}}{\partial P_{i}}P_{i}^{\operatorname{T}}+P_{i}\left[\frac{\partial\operatorname{tr}\big{(}\big{[}P_{i}^{\operatorname{T}}A_{i}P_{i}\big{]}^{2}\big{)}}{\partial P_{i}}\right]^{\operatorname{T}}\right)$
[…]
assuming $\phi_{i}\geq 0$ for $1\leq i\leq\ell$ but no requirement to impose on $\phi_{\ell+i}$ for $1\leq i\leq t$ is necessary, because all $\operatorname{tr}\big{(}\big{[}P_{i}^{\operatorname{T}}A_{i}P_{i}\big{]}^{2}\big{)}$ share the same $\alpha=1$ and $\beta=3$ in (5.5b) while all $\operatorname{tr}(P^{\operatorname{T}}B_{i}P)$ share the same $\underaccent{\bar}{\alpha}=2$ and $\underaccent{\bar}{\beta}=0$ in (8.16b) which is also an equality for the case.
@@
which is positive semidefinite for $x>0$ and $y\geq 0$ if and only if $20\leq\theta\leq 1/$ 2. This implies $\phi(\boldsymbol{x})\equiv\phi_{0}(x_{1},x_{2}+x_{3})$ is convex for $x_{1}>0$ and $x_{2}+x_{3}\geq 0$, provided $20\leq\theta\leq 1/$ 2. However, the results in [67] says that the NEPv approach works on $f$ for $\Theta$TR for $0\leq\theta\leq 1$, much bigger range for $\theta$ than what Theorem 8.2 for $f^{2}$ implies. Theorem 8.2 also yields an inequality on how much the objective function squared, $f^{2}$, increases, in contrast to the previous (6.11) which is for the original objective function, $f$, for $0\leq\theta\leq 1$ and OCCA. In any case, the best $Q_{i}$ at Line 4 of Algorithm 1 when applied to $\Theta$TR is an orthogonal polar factor of $[\widehat{P}^{(i)}]^{\operatorname{T}}D$ and $\mathbb{P}={\rm St}(k,n)_{D+}$.
@@ 9 A Brief Comparison of the NPDo and NEPv Approaches
The NEPv approach requires weaker conditions. We have demonstrated that the NDPo Ansatz demands more on an objective function than the NEPv Ansatz, and hence the NEPv approach provably works on a wider collection of maximization problems (1.1) on the Stiefel manifold than the NPDo approach.
@@
    As a further demonstration, in Table 6, we summarize what are required by both approaches on four convex compositions of matrix-trace functions, and it clearly indicates that NEPv requires weaker conditions than NPDo does for $\phi\circ T_{1a}$, $\phi\circ T_{2a}$, and $\phi\circ T_{4}$.
The NPDo approach is easier to use. The NPDo approach, if it provably works, is easier to implement and more flexible to use.
@@ 10 Concluding Remarks
This is an $n\times k$ matrix equation in $P$ on the Stiefel manifold, upon noticing $\Lambda=[P^{\operatorname{T}}\mathscr{H}(P)+\mathscr{H}(P)^{\operatorname{T}}P]/2$. Any maximizer is a solution. Except for very special objective functions such as $\operatorname{tr}(P^{\operatorname{T}}AP)$ or $\operatorname{tr}(P^{\operatorname{T}}D)$, solving this equation rightly and efficiently for a maximizer is a challenging task. For example, it often has infinitely many solutions and maximizers hide among them. Hence we need to not only solve the nonlinear equation of the KKT condition but also find the right ones. Inspired by recent works [67, 66, 74, 75, 76, 77], in this paper, we establish two unifying frameworks, one for the NEPv approach and the other for the NPDo approach, for optimization on the Stiefel manifold. Our frameworks are built upon two fundamental ansatzes, the NPDo Ansatz and the NEPv Ansatz. When a respective ansatz is satisfied, the corresponding approach, the NPDo or NEPv approach, is guaranteed to work in the sense of global convergence from any given initial point. To expand the applicability of the approaches, we propose the theories of atomic functions for each approach and show that any convex composition $\phi\circ T$ of atomic functions for any of the two approaches satisfies the corresponding ansatz under some mild conditions. It is demonstrated that the commonly used matrix-trace functions
@@
However when an ansatz fails to hold for a given objective function, the conclusions from our global convergence analysis will likely fail. In the case for the NEPv approach, a remedy via the level-shifting SCF exists to ensure locally linear convergence when $f(P)$ is right-unitarily invariant [4] or when $f(P)$ contains and increases with $\operatorname{tr}(P^{\operatorname{T}}D)$ [46], where sharp estimations of linear convergence rate are obtained. But there are more works to do, especially for the NPDo approach for which a remedy remains to be found. In [62], it is investigated how $\operatorname{tr}(P^{\operatorname{T}}D)$ may help determine a particular $P$ among all orthonormal basis matrices of a subspace.
Not all objective functions (or their simple transformations like $f^{2}$ for $\Theta$TR as explained in Remark 8.1) that satisfy the ansatz(es) take the form of some convex compositions of atomic functions, even for some of those in Table 1. For example, in the case of OLDA, $f(P)=\phi\circ T(P)$ where $\phi(\boldsymbol{x})=x_{1}/x_{2}$ and $T(P)=\begin{bmatrix}\operatorname{tr}(P^{\operatorname{T}}AP)\\ \operatorname{tr}(P^{\operatorname{T}}BP)\end{bmatrix}$. Although this $f(P)$ is not a convex composition of $\operatorname{tr}(P^{\operatorname{T}}AP)$ and $\operatorname{tr}(P^{\operatorname{T}}BP)$, it still satisfies the NEPv Ansatz. In fact, OLDA is a special case of $\Theta$TR: $\theta=1$ and $D=0$, and inequality (6.11) applies and yields, for OLDA,